Consider the following information on Budget Plc:
Debt: 80,000 9 coupon bonds outstanding with par value of $1,000 and 18 years to maturity, selling for 108 percent of par, the bonds make semiannual payments.

Common stock: 415,000 shares outstanding, selling for $65 per share: the beta is 1.25

Preferred stock: 100,000 shares of 4.5 percent preferred stock outstanding, currently selling for $103 per share (par value=100)

Market: 8 percent market risk premium and 2.8 percent risk free rate.

Assume the company's tax rate is 35 percent.

1) Calculate the market value of each component of the capital structure (Debt, common stock, and preferred sfock) as well as the total market value of the firm.

2) Calculate the cost of equity.

3) Calculate the after tax cost of debt.

4) Calculate the cost of preferred stock.

5) Calculate the weighted average cost of capital (WACC)
Answer #1
1) MV of equity=Price of equity*number of shares outstanding
MV of equity=65*415000
=26975000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*80000*1.08
=86400000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=103*100000
=10300000
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=26975000+86400000+10300000
=123675000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 26975000/123675000
W(E)=0.2181
Weight of debt = MV of Bond/MV of firm
Weight of debt = 86400000/123675000
W(D)=0.6986
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 10300000/123675000
W(PE)=0.0833
2) Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 2.8 + 1.25 * (8)
Cost of equity% = 12.8
3)Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =18x2
1080 =∑ [(9*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^18x2
                   k=1
YTM = 8.1452785492
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 8.1452785492*(1-0.35)
= 5.29443105698
4) cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 4.5/(103)*100
=4.37
5) WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=5.29*0.6986+12.8*0.2181+4.37*0.0833
WACC =6.85%